# Reformulation of Einstein equations in the Fully Constrained Formulation: local-uniqueness, post-Newtonian expansion and initial data

### v1

@@ III Reformulation of the equations
$N=1-\frac{U}{c^{2}}+\mathcal{O}\left(\frac{1}{c^{4}}\right),\;\psi=1+\frac{U}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{4}}\right),$
$N\psi^{2}=1+\mathcal{O}\left(\frac{1}{c^{4}}\right),$
$\beta^{i}=\mathcal{O}\left(\frac{1}{c^{3}}\right),$
$h^{ij}=\mathcal{O}\left(\frac{1}{c^{4}}\right),$
$\hat{A}^{ij}=\mathcal{O}\left(\frac{1}{c^{3}}\right),\;X^{i}=\mathcal{O}\left(\frac{1}{c^{3}}\right),$
$\hat{A}^{ij}_{TT}=\mathcal{O}\left(\frac{1}{c^{5}}\right),$
@@
We introduce now a new vector field, with the goal of simplifying, at least partially, our equations:
@@
The post-Newtonian order of this vector is $V^{i}=\mathcal{O}(c^{-5})$. This order is justified later in this section. Once $X^{i}$, $\psi$, $N$ and $V^{i}$ are known, the shift vector $\beta^{i}$ can be directly computed from the previous definition for $V^{i}$ (no need to solve the elliptical equation for $\beta^{i}$ anymore). The equation most positively affected by this new auxiliary variable is the evolution equation for the $h^{ij}$ tensor, which, in terms of the new variables, reads:
@@
It turns out that, after the introduction of $V^{i}$, this equation has been slightly simplified and, more importantly, has balanced post-Newtonian orders in both sides of the equation, which are now both $\mathcal{O}(c^{-5})$. The previous version of this equation, Eq. (14), had $\mathcal{O}(c^{-5})$ at the left hand side, while the individual terms on the right hand side were $\mathcal{O}(c^{-3})$. Of course, in the previous version, terms of order $\mathcal{O}(c^{-3})$ must theoretically cancel out to become $\mathcal{O}(c^{-5})$, but this may not always happen numerically. With the introduction of the new vector $V^{i}$, we avoid this potential numerical issue.
Let us now focus on the other evolution equation for $\hat{A}^{ij}$ in Cordero-Carrión et al. (2009). Instead of $\hat{A}^{ij}$, we propose explicitly working with the variables $X^{i}$ and $\hat{A}^{ij}_{TT}$. Notice that $\hat{A}^{ij}$ is $\mathcal{O}(c^{-3})$, while $\hat{A}^{ij}_{TT}$ is $\mathcal{O}(c^{-5})$ in post-Newtonian order. The main reason to take into account $\hat{A}^{ij}_{TT}$ is to solve the spacetime metric more accurately in those cases or regions where the post-Newtonian approximation is valid. In other cases, our reformulation is just a change of variables with no additional effects, beyond the slight reducing of the number of source terms in Eq. (17). The resulting evolution equation for $\hat{A}^{ij}_{TT}$ reads:
@@
where $\dot{X}^{i}=\partial_{t}X^{i}$,
@@
$+\frac{1}{3}\tilde{\gamma}^{ik}X^{j}\mathcal{D}_{k}\mathcal{D}_{j}(2N\psi^{-6})+\mathcal{D}_{k}(2N\psi^{-6})\left(2\tilde{\gamma}^{kj}\mathcal{D}_{j}X^{i}+\frac{1}{3}\tilde{\gamma}^{ik}\mathcal{D}_{j}X^{j}+\frac{1}{3}\tilde{\gamma}^{ij}\mathcal{D}_{j}X^{k}-\hat{A}^{ik}\right)=\mathcal{O}\left(\frac{1}{c^{5}}\right);$
@@ IV.3 Comparison with results in the xCFC formulation
By xCFC formulation we mean imposing CFC condition with the decomposition considered in Eq. (12) and modifications presented in Cordero-Carrión et al. (2009) (where the auxiliary variables $V^{i}$ and $\dot{X}^{i}$ were not introduced). In this subsection we impose $h^{ij}=0$, so we can compare our results with the ones obtained with this xCFC formulation, that can be seen as an approximation to the FCF, except in the spherically symmetric case, where the xCFC formulation is exact (see Cordero-Carrion et al. (2011)). Then, we will compare the numerical solution of Eqs. (21)–(25) imposing $h^{ij}=0$ to the resulting one using the original xCFC formulation. Indeed, it is convenient to solve an elliptic equation for $(N\psi^{2}-1)$ instead of $N\psi^{2}$, because $(N\psi^{2}-1)=\mathcal{O}(c^{-4})$ in PNE. Therefore, the elliptic equations for $X^{i}$ and $\psi$ are the same as in the xCFC case, and the main differences in our proposal are: (i) we compute $N$ from the elliptic equation for $(N\psi^{2}-1)$, derived from Eq. (23); (ii) the shift vector is obtained once $V^{i}$ is known; and (iii) the elliptic equation for $\dot{X}^{i}$ was not considered in the original xCFC formulation.
Figure 1: Radial profiles of $\psi$ and $N$ at $\theta=\pi/2$. The location of the surface of the star at $\theta=\pi/2$ is denoted with a solid black line.
In Figure 1 we show the profiles of the numerical solutions of the variables $\psi$ and $N$. To get $N$, we use the numerical solutions of the elliptic equations for $(N\psi^{2}-1)$ and $\psi$. In Figure 2 we check that the radial profile of $(N\psi^{2}-1)\sim\mathcal{O}(c^{-4})$ decreases much faster than $(N-1)\sim\mathcal{O}(c^{-2})$ and $(\psi-1)\sim\mathcal{O}(c^{-2})$; for example, at $r=10^{2}$, $(N\psi^{2}-1)$ is 100 times smaller than the other two quantities.
@@
Figure 3 shows the numerical solution of the new introduced vector field $V^{i}$ and the shift vector $\beta^{i}$, directly computed from $V^{i}$. For these vector fields only the angular components $V^{\phi}$ and $\beta^{\phi}$ are non-zero. $V^{i}\sim\mathcal{O}(c^{-5})$, while $X^{i}\sim\mathcal{O}(c^{-3})$ and $\beta^{i}\sim\mathcal{O}(c^{-3})$. Radial profiles of these quantities are displayed on Figure 4. When we compare $\beta^{i}$ computed through the xCFC formulation, $\beta_{\text{xCFC}}^{i}$, to that computed with our new approach, the difference is very small (i.e., around $10^{-6}$ at most) and increases with $r$, as it can be observed on Figure 5, where the radial profiles of this difference in absolute value at $\theta=0,\pi/4,\pi/2$ are displayed. This small difference may be due to the fact that we are solving $V^{i}\sim\mathcal{O}(c^{-5})$ (with an elliptic equation with source terms of the same post-Newtonian orders), and then we compute $\beta^{i}$ from $V^{i}$; this strategy may be more accurate for larger radial values with respect to the direct computation of $\beta_{\text{xCFC}}^{i}$.
Figure 3: Radial profiles of $\beta^{\phi}$ and $V^{\phi}$ at $\theta=0$ and $\theta=\pi/2$. The location of the surface of the star at $\theta=0$ and $\theta=\pi/2$ is denoted with dashed and solid black lines, respectively.
@@ IV.5 Convergence
We will compare with more detail our results with those computed with LORENE using several resolutions and placing the outer boundary at different radii. LORENE uses spectral methods and an exponential convergence with increasing resolution for smooth variables is expected. In our case, we use finite differences of second-order, so our convergence with increasing resolution is expected to be slower. Then, we use LORENE solutions as reference for testing the convergence and reliability of our approach. We compute the relative error between our numerical solution $f$ and the one derived with LORENE $f_{L}$,
$\varepsilon_{r}(f)=\frac{||f-f_{L}||_{2}}{||f_{L}||_{2}},$
both with the same spatial resolution and numerical domain. We use the discrete 2-norm,
$||f||_{2}=\left(\displaystyle{\sum_{i,j=1}^{\begin{subarray}{c}i=n_{r},\\ j=n_{\theta}\end{subarray}}}f(r_{i},\theta_{j})^{2}/n_{r}n_{\theta}\right)^{1/2}.$
@@
Regarding the location of the outer boundary, we see a similar behavior with other tested resolutions: $\Delta r=772$ m, $193$ m, $96.4$ m and $48.2$ m. Since we concluded that placing the outer boundary at $r=3R_{\text{star}}$ is enough for guaranteeing the numerical solution not to be strongly affected by the boundary condition, we fix the outer boundary at $r=6R_{\text{star}}$. In Figure 12 we display the relative errors for several spatial resolutions. All components stabilize its relative error for a spatial resolution of about $\Delta r\gtrsim 100$ m.
@@
Figure 12: Relative errors comparing with LORENE solutions for the non-zero components of $h^{ij}$ for several radial spatial resolutions.
@@
$\varepsilon(\Delta r)=|u(\Delta r)-u\left(\Delta r_{\text{hi}}\right)|,$
where $u(\Delta r)$ is the numerical value of the variable $u$ at a fixed point $(r,\theta)$ when a spatial resolution $\Delta r$ is used, and $\Delta r_{\text{hi}}=48.2$ m corresponds to the highest resolution. In Figure 13 we plot these quantities for the non-zero components of $h^{ij}$ with $\Delta r=772/2^{k}$ m, $k=0,1,2,3$, at $\theta=\pi/2$ and $r=0$ km, $r=5$ km. In Table 3 we estimate the order of convergence $p$ at $r=0$ km and $r=5$ km, with $\Delta r=386$ m, according to the following expression
$p\approx\log_{2}\left(\frac{|u(\Delta r)-u(\Delta r/2)|}{|u(\Delta r/2)-u(\Delta r/4)|}\right).$ (45)
$r$ [km] $p(h^{rr})$ $p(h^{\theta\theta})$ $p(h^{\phi\phi})$ $p(h^{r\theta})$
0 1.9654 2.0646 1.9468 2.2798
5 1.9236 2.1470 2.0655 2.3158
Table 3: Estimation of the order of convergence $p$ with the non-zero components of $h^{ij}$ using a spatial resolution $\Delta r=386$ m in expression (45) at $r=0$ km and $r=5$ km.

### v2

@@ III Reformulation of the equations
$\begin{gathered}N=1-\frac{U}{c^{2}}+\mathcal{O}\left(\frac{1}{c^{4}}\right),\;\psi=1+\frac{U}{2c^{2}}+\mathcal{O}\left(\frac{1}{c^{4}}\right),\\ N\psi^{2}=1+\mathcal{O}\left(\frac{1}{c^{4}}\right),\\ \beta^{i}=\mathcal{O}\left(\frac{1}{c^{3}}\right),\\ h^{ij}=\mathcal{O}\left(\frac{1}{c^{4}}\right),\\ \hat{A}^{ij}=\mathcal{O}\left(\frac{1}{c^{3}}\right),\;X^{i}=\mathcal{O}\left(\frac{1}{c^{3}}\right),\\ \hat{A}^{ij}_{TT}=\mathcal{O}\left(\frac{1}{c^{5}}\right),\end{gathered}$ (16)
@@
In light of the previous considerations regarding post-Newtonian orders (16), we reformulate the earlier version of the FCF formulation presented in Cordero-Carrión et al. (2009), aiming to improve numerical accuracy.
For example, Equation (14) motivates the introduction of a new variable involving $\beta^{i}$ and $X^{i}$, both of order $\mathcal{O}(c^{-3})$. $(L\beta)^{ij}$ is the leading term in the first source terms of Equation (14), while $2N\psi^{-6}(LX)^{ij}$ is the leading term in the last source terms of Equation (14). Accordingly, we introduce a new vector field with the goal of simplifying, at least partially, this equation:
@@
It turns out that the post-Newtonian order of this vector is $V^{i}=\mathcal{O}(c^{-5})$, as it will be justified later in this section. Dealing with $V^{i}$ instead of $\beta^{i}$ we expect to solve the spacetime metric more accurately in those cases or regions where the post-Newtonian approximation is valid. Once $X^{i}$, $\psi$, $N$ and $V^{i}$ are known, the shift vector $\beta^{i}$ can be directly computed from the previous definition for $V^{i}$ (no need to solve the elliptical equation for $\beta^{i}$ anymore). The equation most positively affected by this new auxiliary variable is the evolution equation for the $h^{ij}$ tensor, which, in terms of the new variables, reads:
@@
After the introduction of $V^{i}$, this equation has been slightly simplified and, more importantly, has balanced post-Newtonian orders in both sides of the equation, which are now both $\mathcal{O}(c^{-5})$. The previous version of this equation, Eq. (14), had $\mathcal{O}(c^{-5})$ at the left hand side, while the individual terms on the right hand side were $\mathcal{O}(c^{-3})$. Of course, in the previous version, terms of order $\mathcal{O}(c^{-3})$ must theoretically cancel out to become $\mathcal{O}(c^{-5})$, but this may not always happen numerically. With the introduction of the new vector $V^{i}$, we avoid this potential numerical issue.
Let us now focus on the other evolution equation for $\hat{A}^{ij}$ in Cordero-Carrión et al. (2009). Instead of $\hat{A}^{ij}$, we propose explicitly working with the variables $X^{i}$ and $\hat{A}^{ij}_{TT}$. Notice that $\hat{A}^{ij}$ is $\mathcal{O}(c^{-3})$, while $\hat{A}^{ij}_{TT}$ is $\mathcal{O}(c^{-5})$ in post-Newtonian order. Therefore, taking $\hat{A}^{ij}_{TT}$ as variable will solve the spacetime metric more accurately too when the post-Newtonian approximation is valid. In other cases, our reformulation is just a change of variables with no additional effects, beyond the slight reducing of the number of source terms in Eq. (18). The resulting evolution equation for $\hat{A}^{ij}_{TT}$ reads:
@@
where we have introduced the new variable $\dot{X}^{i}=\partial_{t}X^{i}$ (for which we derive an elliptic equation below),
@@
$+\frac{1}{3}\tilde{\gamma}^{ik}X^{j}\mathcal{D}_{k}\mathcal{D}_{j}(2N\psi^{-6})+\mathcal{D}_{k}(2N\psi^{-6})\left(2\tilde{\gamma}^{kj}\mathcal{D}_{j}X^{i}+\frac{1}{3}\tilde{\gamma}^{ik}\mathcal{D}_{j}X^{j}+\frac{1}{3}\tilde{\gamma}^{ij}\mathcal{D}_{j}X^{k}-\hat{A}^{ik}\right)=\mathcal{O}\left(\frac{1}{c^{5}}\right).$
In the final equation (25), we clearly observe (by referring to (16)) what was previously stated regarding the post-Newtonian order of $V^{i}$. Notice also the improvement in numerical accuracy dealing with the variable $N\psi^{2}$, for which we have the elliptic equation (24), instead of $N\psi$ as it was proposed in Cordero-Carrión et al. (2009).
Finally, by taking the divergence of Equation (19), we obtain an elliptic equation for the new variable $\dot{X}^{i}$:
@@ IV.3 Comparison with results in the xCFC formulation
By xCFC formulation we mean imposing CFC condition with the decomposition considered in Eq. (12) and modifications presented in Cordero-Carrión et al. (2009) (where the auxiliary variables $V^{i}$ and $\dot{X}^{i}$ were not introduced). In this subsection we impose $h^{ij}=0$, so we can compare our results with the ones obtained with this xCFC formulation, that can be seen as an approximation to the FCF, except in the spherically symmetric case, where the xCFC formulation is exact (see Cordero-Carrion et al. (2011)). Then, we will compare the numerical solution of Eqs. (22)–(26) imposing $h^{ij}=0$ to the resulting one using the original xCFC formulation. Indeed, it is convenient to solve an elliptic equation for $(N\psi^{2}-1)$ instead of $N\psi^{2}$, because $(N\psi^{2}-1)=\mathcal{O}(c^{-4})$ in PNE. Therefore, the elliptic equations for $X^{i}$ and $\psi$ are the same as in the xCFC case, and the main differences in our proposal are: (i) we compute $N$ from the elliptic equation for $(N\psi^{2}-1)$, derived from Eq. (24); (ii) the shift vector is obtained once $V^{i}$ is known; and (iii) the elliptic equation for $\dot{X}^{i}$ was not considered in the original xCFC formulation. In Figure 1 we check that the radial profile of $(N\psi^{2}-1)\sim\mathcal{O}(c^{-4})$ decreases much faster than $(N-1)\sim\mathcal{O}(c^{-2})$ and $(\psi-1)\sim\mathcal{O}(c^{-2})$; for example, at $r=10^{2}$, $(N\psi^{2}-1)$ is 100 times smaller than the other two quantities.
@@
Radial profiles of $V^{i}$ and the shift vector $\beta^{i}$ (directly computed from $V^{i}$) are displayed on Figure 2. For these vector fields only the angular components $V^{\phi}$ and $\beta^{\phi}$ are non-zero. Let us remember that $V^{i}\sim\mathcal{O}(c^{-5})$, while $X^{i}\sim\mathcal{O}(c^{-3})$ and $\beta^{i}\sim\mathcal{O}(c^{-3})$. We check in Figure 2 that $V^{i}$ decreases much faster than $X^{i}$ and $\beta^{i}$ getting to differences of almost two orders of magnitude at the outer boundary. When we compare $\beta^{i}$ computed through the xCFC formulation, $\beta_{\text{xCFC}}^{i}$, to that computed with our new approach, the difference is very small (i.e., around $10^{-6}$ at most) and increases with $r$, as it can be observed on Figure 3, where the radial profiles of this difference in absolute value at $\theta=0,\pi/4,\pi/2$ are displayed. This small difference may be due to the fact that we are solving $V^{i}\sim\mathcal{O}(c^{-5})$ (with an elliptic equation with source terms of the same post-Newtonian orders), and then we compute $\beta^{i}$ from $V^{i}$; this strategy may be more accurate for larger radial values with respect to the direct computation of $\beta_{\text{xCFC}}^{i}$.
@@ IV.5 Convergence
We will compare with more detail our results with those computed with LORENE using several resolutions and placing the outer boundary at different radii. LORENE uses spectral methods and an exponential convergence with increasing resolution for smooth variables is expected. In our case, we use finite differences of second-order, so our convergence with increasing resolution is expected to be slower. Then, we use LORENE solutions as reference for testing the convergence and reliability of our approach. We compute the relative error between our numerical solution $f$ and the one derived with LORENE $f_{L}$, $\varepsilon_{r}$, using the discrete 2-norm.
@@
Regarding the location of the outer boundary, we see a similar behavior with other tested resolutions: $\Delta r=772$ m, $193$ m, $96.4$ m and $48.2$ m. Since we concluded that placing the outer boundary at $r=3R_{\text{star}}$ is enough for guaranteeing the numerical solution not to be strongly affected by the boundary condition, we fix the outer boundary at $r=6R_{\text{star}}$. All components stabilize its relative error for a spatial resolution of about $\Delta r\gtrsim 100$ m.